OpenStudy (anonymous):

Help With Algebra 1!

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system −3 3

OpenStudy (anonymous):

A) x can be any number as there are infinitely many solutions to this system B)There is no x value as there is no solution to this system C) -3 D) 3

iYuko (iyuko):

Sub in "-3" for x

iYuko (iyuko):

in both equations tell me if they equal

OpenStudy (anonymous):

What do you mean @iYuko

OpenStudy (texaschic101):

4x - 3y = 3...(4) 5x - 4y = 3....(-3) ------------ 16x - 12y = 12 -15x + 12y = -9 ---------add x = 3

OpenStudy (anonymous):

Thank you! I'll fan and medal you @texaschic101

OpenStudy (texaschic101):

yw